2. By die Internation meth (10 ) JA
Answer #1

\small Centroid \: on\: \, the\: x-axis,\: \overline{x}=\frac{\int xdA}{\int dA}

yo 1-y=5 (1 da HE- w dA=ydn = (in - x²) da [dA=(5x-x²) d x] > we need to find the intersection of these curves. Oily = sa (i)

- -[-] = } - $ = 5 -> Jmda = f** (v2 – x) dx = [ (054 - 23) Ix - 3 = 5(2%-x²) da OM 3 - = Inda = 3 --07 - = B 3/20 n-Unda -=

\small Centroid \: on\: \, the\: x-axis,\: \overline{x}=0.45